OpenStudy (anonymous):

Write the equation of the line which passes through (0, -4) and is parallel to the line 3x - y = 6.

OpenStudy (anonymous):

First find the y= mx + b format of this equation 3x - y = 6 -y = -3x + 6 y = 3x - 6 Parallel equations have the same slope In this case, the slope is 3, making the slope of the parallel equation also 3. Now plug in your coordinates and slope into this equation (y-y1) = m (x - x1) (0,-4) = (x1,y1) m = 3

OpenStudy (anonymous):

Perfect

OpenStudy (anonymous):

thankx

OpenStudy (anonymous):

These are my choices y = 3x - 1 y = 3x + 4 y = 3x - 4 y = 3x + 1 I picked y = 3x - 1

OpenStudy (anonymous):

Its y = 3x - 4 When you plug all the values in: ( y- y1) = m (x-x1) = (y -(-4) = 3 (x -0) = ( y + 4) = 3(x-0) =y + 4 = 3x - 0 y = 3x -4

OpenStudy (anonymous):

3x-y-4=0